OpenStudy (anonymous):

f it is equally likely for a paintball to hit any spot on the practice wall, what is the probability that it will shoot a paintball into section 4? there are four sections ,3 and 4 on the bottom are bigger though so thats what confuses me but it says equally likely so my guess is 1/4

OpenStudy (anonymous):

cant really do this with out the picture. can you up load it? the answer will be the ratio of the area of section 4 to the total area

OpenStudy (anonymous):



OpenStudy (anonymous):

damn a geometry problem it is definitely not \(\frac{1}{4}\) but we need another piece of information, a measurement of some kind, like how high up is the line through the triangle?

OpenStudy (anonymous):

The height of each individual numbered section is 38 inches.

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

you got answers now here http://openstudy.com/study#/updates/4fbc518de4b0556534315e72

OpenStudy (anonymous):

i got 3/8

OpenStudy (zarkon):

I got 3/8 too

OpenStudy (anonymous):

my geometry sucks why is it \(\frac{3}{8}\) ?

OpenStudy (zarkon):

 area of (1) and (2) is \(\frac{1}{2}xa\) each area of (3) and (4) is \(\frac{1}{2}2x2a-\frac{1}{2}xa=\frac{3}{2}xa\)